How do you get a perimeter of 21 with and area of 5 units squared thank you :D
Answer by scott8148(6628)

2x + 2y = 21

xy = 5 ___ x = 5/y

2(5/y) + 2y = 21

multiplying by y ___ 10 + 2y^2 = 21y ___ 2y^2 - 21y + 10 = 0

factoring ___ (2y - 1)(y - 10) = 0

1/2 and 10 are the width and length